# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ 5 Simulation studies
In this section, we assess the relative performances of the proposed estimation algorithms compared to state-of-the-art methods through several simulation experiments. In particular, we are interested in evaluating the considered approaches in terms of execution time, out-of-sample prediction error, and biological signal extraction quality.
@@ 5.1 Data generating process
In our experiments, we considered the following simulation setting: each dataset contains cells from five well-separated types evenly distributed in the sample. The data are also divided into three batches having different expression levels. No lineage or branching effects are considered.
To evaluate the performance of the proposed method under different regimes, we consider two simulation settings. In simulation setting A, we compare several matrix factorization models and algorithms under a fixed latent space rank, $d=5$, and we let the dimensions of the response matrix increase. Specifically, we set the number of cells, $n$, to be 10 times the number of genes, $m$, and we set $m\in\{100,250,500,750,1000\}$. In simulation setting B, we compare the same set of factorization methods by fixing the dimensions of the response matrix to $n=5000$ and $m=500$, and letting the latent space rank grow, i.e., $d\in\{5,10,15,20,25\}$. For each combination of latent rank, $d$, number of cells, $n$, and number of genes, $m$, under the two scenarios, we generated 100 expression matrices.
Additional details are provided in the supplementary material. Moreover, the code to generate the data and run the simulations is publically available on GitHub [^2] Refer to the GitHub repository github/alexandresegers/sgdGMF_Paper.
@@ 5.2 Competing methods and performance measures
  • AIRWLS: Poisson GMF model estimated using the alternated iterative re-weighted least squares algorithm of Kidziński et al., (2022) and Wang and Carvalho, (2023) (sgdGMF package; this work);
@@
  • Newton: Poisson GMF model estimated using the exact quasi-Newton algorithm of Kidziński et al., (2022) described in Algorithm 1 (sgdGMF package; this work);
@@
  • SGD: Poisson GMF model estimated using the proposed adaptive stochastic gradient descent with block-wise subsampling described in Algorithm 2 (sgdGMF package; this work)
@@
Some of the most relevant characterizing features of all these methods and the relative implementations are summarized in Table 1. To emulate a conventional usage of all these packages, we tried to adhere closely to the standard option setups recommended in the documentation provided by their respective authors. All the algorithms implemented in the sgdGMF package are initialized with Algorithm 4. For more details, we refer to the Appendix. Additionally, for all methods allowing parallel computing, we run the estimation employing 4 cores. Notice that such a configuration is commonly supported by modern PCs.
@@
To assess the relative performance of the models under consideration, we train them on a designated training set and subsequently evaluate their goodness-of-fit using a validation set. In each simulation scenario, we initially generate a complete data matrix and then construct the training set by introducing a predetermined percentage of missing values, typically set at $30\%$. The positions of such holdout entries are sampled from a uniform distribution on the matrix indices. The corresponding test set comprises all observations withheld during the training phase. We evaluate the models in terms of elapsed execution time (in seconds), out-of-sample reconstruction, and cell-type separation in the estimated latent space. Let $\mathcal{T}$ and $\mathcal{V}$ be the index sets corresponding to the training and validation entries of the response matrix and let $\bar{y}_{{}_{\mathcal{T}}}$ be the empirical average of the training matrix. The out-of-sample reconstruction error is then computed using the relative Poisson deviance, that is
@@
$\text{Deviance}(\mathbf{Y},\hat{\boldsymbol{\mu}})=\sum_{(i,j)\in\mathcal{V}}D(y_{ij},\hat{\mu}_{ij})\bigg{/}\sum_{(i,j)\in\mathcal{V}}D(y_{ij},\bar{y}_{{}_{\mathcal{T}}}).$
To assess the degree of cell-type cluster separation over the estimated latent space, we compute two validation scores: the average silhouette width (Rousseeuw,, 1987) computed on a two-dimensional t-SNE embedding (Van der Maaten and Hinton,, 2008), and the neighborhood purity (Manning et al.,, 2008) evaluated on the original latent space.
@@ 5.3 Simulation results
Figure 2: Summary information for the simulation experiment described in Section 5.1 with $n=5000$, $m=500$ and $d=5$. Left: summary statistics reporting the execution time (in seconds), the out-of-sample relative Poisson deviance (multiplied by $100\%$), the silhouette score of the true cell-type clusters calculated on a 2-dimensional tSNE projection, and the mean cluster purity of the true cell-type clusters calculated on the 5-dimensional estimated latent space. Right: 2-dimensional tSNE projections of the estimated latent factors for one specific replication of the experiment.
@@
Fig. 2 presents an overview of the results obtained from the data simulated with $n=5000$, $m=500$, and $d=5$. Panel A shows that the proposed SGD method is the fastest implementation, while performing similarly to the best performing methods in terms of out-of-sample error and silhouette width. Panel B shows, on an exemplar simulation run, that all the GMF methods are able to reconstruct the original cell-type clustering, with the exception of AvaGrad, which is unable to separate the different groups. On the other hand the NMF methods over-cluster the data, unable to remove the simulated batch effects. This is confirmed by the average silhouette width (bottom plot of panel A) that shows a large difference in the performance of NMF and GMF methods.
@@
These results are confirmed across both setting A and B (Fig. 3): in terms of computational efficiency, the proposed SGD implementation consistently outperforms its competitors across both settings, showing lower elapsed execution times and a better scalability with respect to both the sample size and the dimension of the latent space (Fig. 3A). Among the alternative methods, Newton emerges as the fastest, while NBWaVE and AIRWLS exhibit similar performances, which is unsurprising given their shared algorithmic strategies. On the other hand, Fisher and AvaGrad optimizers display inferior scalability compared to SGD, Newton, AIRWLS and NBWaVE (Fig. 3A). Contrary to the claims in the glmPCA package documentation, our findings indicate that the AvaGrad optimizer often lacks in both speed and reliability when compared to the Fisher optimizer, which typically achieves convergence within a reasonable time and, on average, runs faster than AvaGrad.
@@
Regarding non-negative matrix factorization, implementations such as NMF and CMF  demonstrate poor computational scalability across both settings. However, NMF+ shows improved performance, reaching efficiency levels close to NBWaVE and AIRWLS (Fig. 3A).
In terms of goodness-of-fit, a notable distinction arises between non-negative and generalized matrix factorization methods. Typically, GMF-based approaches consistently yield lower values of relative residual deviance compared to their NMF-based counterparts (top row of Fig. 3B). However, two exceptions to this trend are evident with NMF+ performing well and AvaGrad performing poorly. Among all compared methods, SGD consistently emerges as the most accurate approach in terms of residual deviance (top row of Fig. 3B).
@@
The difference in performance between NMF and GMF methods becomes even clearer when looking at the silhouette width and neighborhood purity scores (middle and bottom rows of Fig. 3B). In both scenarios A and B, NBWaVE, AIRWLS, and Newton consistently outperform CMF, NMF, and NMF+. While SGD exhibits slightly suboptimal performance compared to NBWaVE, AIRWLS, and Newton, especially in small sample settings, it converges towards them as the sample size increases. This behaviour is not unexpected, as stochastic optimization algorithms typically require a large number of observations to mitigate their intrinsic randomness, achieving stability as the data dimension grows.
@@
Figure 3: Summary statistics of the simulation experiments described in Section 5.1. Left: elapsed execution time (in seconds) under the simulation settings A and B. Right: goodness of fit measures under the simulation settings A (left column) and B (right column). The considered goodness of fit measures are: the out-of-sample normalized residual deviance, the silhouette evaluated on a 2-dimensional tSNE protection of the latent space, the neighborhood purity of the true cell-type evaluated on the original latent space.
@@ 6 Real data applications
Further, we apply our method to a large scRNA-seq dataset consisting of more than 1.3 million cells from the mouse brain, generated by 10X Genomics (Lun and Morgan,, 2023). This dataset showcases the scalability of our approach in large datasets.
Throughout this section, we consider parametrization (B) to obtain an orthonormal loading matrix $\mathbf{V}$ and a scaled orthogonal score factor matrix $\mathbf{U}$. This choice is conventional in the RNA–seq literature and is coherent with the standard parametrization of principal component analysis.
@@ 6.1 Arigoni data
The original Arigoni dataset (Arigoni et al.,, 2024) consist of $29{,}606$ cells from 8 different cell lines. Quality control is done using the perCellQCFilters function of the R package scuttle (McCarthy et al.,, 2017), which filtered cells that have a library size lower than 1306, a percentage of mitochondrial reads higher than 6.05% or cells that have fewer than 732 features expressed. Additionally, peripheral blood mononuclear cells are removed due to their distinct expression profile. The final filtered dataset includes $26{,}472$ cells from 7 different cell lines. Unless mentioned differently, all the analyses are based on the 500 most variable genes, a common choice in standard scRNA-seq workflows.
@@
To select the optimal matrix rank for the latent component of the model, different model selection criteria are assessed in a 5-fold cross validation, assigning missing values to the test partition of each fold. The model selection criteria are shown in Panel A of Figure 4, with Panel B and C showing the evaluation of the model selection criteria. AIC and BIC are assessed on the 5 training data partitions, while the out-of-sample deviance is calculated on the test data. Also, considering all the data, we assess the scree plot of the eigenvalues based on the deviance residuals after using OLS on the log-transformed data.
@@
Both the AIC and out-of-sample deviance criteria suggest a matrix rank of 15, while the BIC and the scree plot suggest 9 (Fig. 4A). The mean cell-line neighborhood purity scores (Fig. 4B) show that, for the majority of cell lines, a rank of 9 is sufficient to completely separate the groups, with only cell lines A549 and CCL-185-IG exhibiting a score lower than $0.9$. At rank 15, all cell lines achieve a large value of mean purity and increasing the rank further does not improve this index, while introducing the risk of overfitting the data, as observed in the increasing out-of-sample deviance in the cross validation (Fig. 4A). These remarks are confirmed when visually inspecting the tSNE plots, coloured by the ground truth labels (Fig. 4C), that show that a matrix rank of 9 yields a good separation, except for A549 from CCL-185-IG, which show a slight degree of mixing (Fig. 4C, confusion matrix). Although there are no outstanding visual differences between rank 15 and 30, performing Leiden clustering with a resolution tuned to obtain 7 clusters shows that a matrix of rank 30 results in one small cluster with very few cells, rather than separating the A549 and CCL-185-IG groups, while this happens as expected for rank 15. This suggests that 15 is a reasonable number of latent factors to include in the model.
@@
Using 15 factors, as suggested by the model selection criteria, SGD was compared with NBWaVE, Fisher and AvaGrad, three of the most popular methods in single-cell analysis that perform well in the simulations (Supplementary Fig. 1). All three methods achieve similar results, as shown by both visual inspection of the tSNE plots and the mean cell-line neighborhood purities. However, SGD is orders of magnitude faster, and therefore allows for model selection obtaining optimal matrix rank, which was shown to be important for clustering the data. Further, SGD has a better out-of-sample deviance, probably because it can deal with missing values, unlike NBWaVE and Fisher and AvaGrad which require imputation.

### v2

@@ 5 Simulation studies
In this section, we assess the relative performances of the proposed estimation algorithms compared to state-of-the-art methods through several simulation experiments. In particular, we are interested in evaluating the considered approaches in terms of execution time, memory consumption, out-of-sample prediction error, and biological signal extraction quality.
@@ 5.1 Data generating process
In our experiments, we considered the following simulation setting: each dataset contains cells from five well-separated types evenly distributed in the sample. This is the signal that we aim to reconstruct with the latent factors. Moreover, the data are divided into three batches having different expression levels. This effect can be modeled as row covariates in our approach. No lineage or branching effects are considered. Moreover, the simulation includes cell-specific library sizes, which mimic the differences in the total number of counts per cell observed in real data. We model such effects with a column intercept in our framework.
To evaluate the performance of the proposed method under different regimes, we consider two simulation settings. In simulation setting A, we compare several matrix factorization models and algorithms under a fixed latent space rank, $d=5$, and we let the dimensions of the response matrix increase. Specifically, we set the number of cells, $n$, to be 10 times the number of genes, $m$, and we set $m\in\{100,250,500,750,1000\}$. In simulation setting B, we compare the same set of factorization methods by fixing the dimensions of the response matrix to $n=5000$ and $m=500$, and letting the latent space rank grow, i.e., $d\in\{5,10,15,20,25\}$. For each combination of latent rank, $d$, number of cells, $n$, and number of genes, $m$, under the two scenarios, we generated 100 expression matrices. Additional details are provided in Appendix C in the Supplementary Material. Moreover, the code to generate the data and run the simulations is publicly available on GitHub [^2] Refer to the GitHub repository github/alexandresegers/sgdGMF_Paper.
@@ 5.2 Competing methods and performance measures
GFM-AM: Poisson generalized factor model estimated via alternated maximization (GFM package; Liu et al.,, 2023);
GFM-VEM: Poisson generalized factor model estimated via variational expectation maximization (GFM package; Nie et al.,, 2024);
COAP: Covariate-augmented overdispersed Poisson factor model via variational expectation maximization (COAP package; Liu and Zhong,, 2024);
@@
AIRWLS: Poisson GMF model estimated via the alternated iterative re-weighted least squares algorithm of Kidziński et al., (2022) and Wang and Carvalho, (2023) (sgdGMF package);
@@
Newton: Poisson GMF model estimated via the exact quasi-Newton algorithm of Kidziński et al., (2022) (sgdGMF package);
@@
aSGD: Poisson GMF model estimated via the proposed adaptive stochastic gradient descent with block-wise subsampling described in Algorithm 1 (sgdGMF package; this work)
@@
Some of the most relevant characterizing features of all these methods and the relative implementations are summarized in Table 1. To emulate a conventional usage of all these packages, we tried to adhere closely to the standard option setups recommended in the documentation provided by their respective authors. All the algorithms implemented in the sgdGMF package are initialized with the strategy described in Section 4.1. Additionally, for all methods allowing parallel computing (CMF, NMF+, NBWaVE, AIRWLS and Newton), we run the estimation employing 4 cores, a configuration commonly supported by modern PCs. For more details, we refer to Appendix C in the Supplementary Material.
To account for technical confounders, we use the batch group as a covariate in all the models supporting regression effects, namely CMF, AvaGrad, Fisher, NBWaVE, COAP, AIRWLS, Newton, and aSGD. Additionally, we also included row- and column-specific intercepts to account for cell- and gene-specific effects. Thus, in our model formulation, we obtain the linear predictor $\eta_{ij}=\gamma_{0i}+\mathbf{x}_{i:}^{\top}\boldsymbol{\beta}_{:j}+\mathbf{u}_{i:}^{\top}\mathbf{v}_{:j}$, where $\mathbf{x}_{i:}$ is a vector of dummy variables identifying which batch each cell $i$ belongs to.
@@
To assess the relative performance of the models under consideration, we estimate them on a designated training set and subsequently evaluate their goodness-of-fit using a validation set. In each simulation scenario, we initially generate a complete data matrix and then construct the training set by introducing a predetermined percentage of missing values, typically set at $30\%$. The positions of such holdout entries are sampled from a uniform distribution on the matrix indices. The corresponding test set comprises all observations withheld during the training phase. We evaluate the models in terms of elapsed execution time (in seconds), peak random access memory consumption (in megabytes), out-of-sample reconstruction, and cell-type separation in the estimated latent space. Let $\mathcal{T}$ and $\mathcal{V}$ be the index sets corresponding to the training and validation entries of the response matrix, and let $\bar{y}_{{}_{\mathcal{T}}}$ be the empirical average of the training matrix. The out-of-sample reconstruction error is then computed using the relative logarithmic root mean squared error, and the relative Poisson deviance, which are defined as
@@
$\text{Error}(\mathbf{Y},\hat{\boldsymbol{\mu}})=\frac{\sum_{(i,j)\in\mathcal{V}}\big[\log\{(1+y_{ij})/(1+\hat{\mu}_{ij})\}\big]^{2}}{\sum_{(i,j)\in\mathcal{V}}\big[\log\{(1+y_{ij})/(1+\bar{y}_{{}_{\mathcal{T}}})\}\big]^{2}},\quad\text{Deviance}(\mathbf{Y},\hat{\boldsymbol{\mu}})=\frac{\sum_{(i,j)\in\mathcal{V}}D(y_{ij},\hat{\mu}_{ij})}{\sum_{(i,j)\in\mathcal{V}}D(y_{ij},\bar{y}_{{}_{\mathcal{T}}})}.$
To assess the degree of cell-type cluster separation over the estimated latent space, we consider two validation scores: the average silhouette width (Rousseeuw,, 1987) computed on a two-dimensional tSNE embedding (Van der Maaten and Hinton,, 2008), and the neighborhood purity (Manning et al.,, 2008) evaluated on the original latent space, which can be evaluated using the functions silhouette() and neighborPurity() from the R packages cluster (Maechler et al.,, 2022) and bluster (Lun,, 2023), respectively. The average silhouette is a global measure of cluster cohesion ranging from 1-1 to $1$, with $1$ indicating perfect separation between clusters. The neighborhood purity is a local measure of cluster cohesion ranging from 0 to $1$, with $1$ indicating perfect coherence between clusters. Being a global measure based on Euclidean distances, the silhouette favors clusters with spherical shapes, while it is not able to detect well-separated clusters featuring non-spherical or non-convex boundaries. On the other hand, neighborhood purity is a local measure of cluster separation, which is able to detect localized behavior and does not depend on the cluster shapes.
@@ 5.3 Simulation results
Figure 2: Summary information for the simulation experiment described in Section 5.1 with $n=5000$, $m=500$ and $d=5$. Left (panel A): summary statistics reporting the execution time (in seconds), the peak memory consumption (in megabytes), the out-of-sample relative logarithmic root mean squared error (multiplied by $100\%$), the out-of-sample relative Poisson deviance (multiplied by $100\%$), the silhouette score of the true cell-type clusters calculated on a 2-dimensional tSNE projection, and the mean cluster purity of the true cell-type clusters calculated on the 5-dimensional estimated latent space. Right (panel B): 2-dimensional tSNE projections of the estimated latent factors for one specific replication of the experiment.
@@
Figure 2 presents an overview of the results obtained from the data simulated with $n=5000$, $m=500$, and $d=5$. Our proposed aSGD method is the fastest approach together with GFM-VEM and COAP (Fig. 2A): it is parsimonious in terms of memory usage, while performing similarly to the best-performing methods in terms of out-of-sample error, deviance, silhouette width, and neighborhood purity. An exemplar simulation run (Fig. 2B) shows that all the GMF methods using covariate information succeed in reconstructing the original cell-type clustering, being able to filter out the batch effect via the regression term in the linear predictor. The only exception is AvaGrad, which fails to separate the different groups. On the other hand, GFM-AM and GFM-VEM are unable to disentangle the cell-types from the batch effects, primarily due to the GFM package’s inability to account for covariate effects. Similarly, the NMF methods over-cluster the data, not allowing for batch effect removal via regression. This is confirmed by the average silhouette width and the mean neighborhood purity (bottom lines of Fig. 2A), which show a large difference in performance, discriminating methods based on their ability to account for covariate (i.e., batch) effects.
@@
These results are confirmed across both setting A and B (Fig. 3 and 4): in terms of computational efficiency, the proposed aSGD implementation consistently outperforms its competitors in setting A and ranks as the top performer in setting B, showing lower elapsed execution times, and a better scalability with respect to both the sample size and the dimension of the latent space (top row of Fig. 3). Additionally, aSGD manifests a parsimonious management of the random access memory, which is aligned to AIRWLS, Newton, and GFM-VEM and surpassed only by CMF and NMF+ (bottom row of Fig. 3). Interestingly, CMF, NMF+, GFM-VEM, AIRWLS, Newton, and aSGD maintain an almost constant memory footprint in simulation setting B. This behavior is attributed to their efficient memory allocation strategies, where memory usage is mainly dictated by the storage of input data and sufficient statistics, with negligible additional cost during optimization, even as the latent dimension scales but is much smaller than the matrix dimensions. Among the alternative methods, COAP and Newton emerge as the fastest, followed by AIRWLS. On the other hand, GFM-AM, NBWaVE, Fisher and AvaGrad optimizers display inferior scalability compared to aSGD, Newton, and AIRWLS (Fig. 3). Contrary to the claims in the glmpca package documentation, our findings indicate that the AvaGrad optimizer often lacks in both speed and reliability when compared to the Fisher optimizer, which typically achieves convergence within a reasonable time and, on average, runs faster than AvaGrad.
@@
Regarding non-negative matrix factorization, implementations such as NMF and CMF demonstrate poor computational scalability across both settings. However, NMF+ shows improved performance, reaching efficiency levels close to AIRWLS in setting B, and an optimal management of the memory in both settings (Fig. 3).
@@
In terms of goodness-of-fit measures (top two rows of Fig. 4), the least out-of-sample logarithmic error and deviance are systematically achieved by NMF+, COAP, AIRWLS, Newton, and aSGD, with aSGD consistently emerging as the most accurate method in the deviance metric. Such a situation partially reflects on the silhouette width and neighborhood purity scores (bottom two rows of Fig. 4), for which, in both scenarios A and B, NBWaVE, COAP, AIRWLS, Newton, and aSGD always outperform the other methods in terms of cell-type separation in the latent space. As such, aSGD exhibits slightly suboptimal performance compared to the other top-performing methods, especially in small sample settings; however, it converges towards them as the sample size increases. This behavior is not unexpected, as stochastic optimization algorithms typically require a large number of observations to mitigate their intrinsic randomness, achieving stability as the data dimension grows.
Figure 3: Summary statistics of the simulation experiments described in Section 5.1. The columns correspond to simulation settings A (left) and B (right). The rows correspond to the elapsed execution time in seconds (top) and the peak memory consumption in megabytes (bottom).
@@
Figure 4: Summary statistics of the simulation experiments described in Section 5.1. The columns correspond to simulation settings A (left) and B (right). The rows correspond to four goodness-of-fit measures. From top to bottom: the out-of-sample relative logarithmic root mean squared error, the out-of-sample relative residual deviance, the silhouette evaluated on a 2-dimensional tSNE projection of the latent space, and the neighborhood purity of the true cell-type evaluated on the original latent space.
@@ 6 Real data applications
Further, we apply our method to a large scRNA-seq dataset consisting of more than 1.3 million cells from the mouse brain, generated by 10X Genomics (Lun and Morgan,, 2023). As this dataset does not contain a ground truth of cell labels, this dataset primarily showcases the scalability of our approach in large datasets.
Throughout this section, we consider parametrization (B1) to obtain an orthonormal loading matrix $\mathbf{V}$ and a scaled orthogonal score factor matrix $\mathbf{U}$ (see Section 2.1). This choice is conventional in the RNA-seq literature (see, e.g., Risso et al.,, 2018; Townes et al.,, 2019; Ahlmann-Eltze and Huber,, 2023) and is coherent with the standard parametrization of principal component analysis.
@@ 6.1 Arigoni data
The original Arigoni dataset (Arigoni et al.,, 2024) consists of 29,60629{,}606 cells from 8 different lung cancer cell lines with unique driver mutations (EGFR, ALK, MET, ERBB2, KRAS, BRAF, ROS1). The ground truth knowledge of the driver mutations can be used to evaluate cell clustering by visual inspection and upon using clustering algorithms. Importantly, the CCL-185-IG cell line is derived from the A549 cell line. Therefore, only subtle differences are expected between these two cell lines. Quality control is done using the perCellQCFilters function of the R package scuttle (McCarthy et al.,, 2017), which filtered cells that have a library size lower than 1306, a percentage of mitochondrial reads higher than 6.05% or cells that have fewer than 732 features expressed. Additionally, peripheral blood mononuclear cells are removed due to their distinct expression profile. The final filtered dataset includes 26,47226{,}472 cells from 7 different cell lines. Unless mentioned differently, all the analyses are based on the 500 most variable genes, a common choice in standard scRNA-seq workflows. The selection of the number of highly-variable genes did not prove critical, as the results are robust across a large spectrum of values (Supplementary Fig. 1 and 2).
@@
To select the optimal matrix rank for the latent component of the model, different model selection criteria are assessed in a 5-fold cross validation (see Section 4.2 for details). AIC and BIC are assessed on the 5 training data partitions, while the out-of-sample deviance is calculated on the test data. Also, considering all the data, we assess the scree plot of the eigenvalues based on the deviance residuals after using OLS on the log-transformed data.
@@
Both the AIC and out-of-sample deviance criteria suggest a matrix rank of 15, while the BIC and the scree plot suggest 9 (Fig. 5A). The mean cell-line neighborhood purity scores (Fig. 5B) show that, for the majority of cell lines, a rank of 9 is sufficient to completely separate the groups, with only cell lines A549 and CCL-185-IG exhibiting a score lower than $0.9$. At rank 15, all cell lines achieve a large value of mean purity, and increasing the rank further does not improve this index, while introducing the risk of overfitting the data, as observed in the increasing out-of-sample deviance in the cross-validation (Fig. 5A). These remarks are confirmed when visually inspecting the tSNE plots, coloured by the ground truth labels (Fig. 5C), that show that a matrix rank of 9 yields a good separation, except for A549 from CCL-185-IG, which show a slight degree of mixing (Fig. 5C, confusion matrix). Although there are no outstanding visual differences between rank 15 and 30, performing Leiden clustering with a resolution tuned to obtain 7 clusters shows that a matrix of rank 30 results in one small cluster with very few cells, rather than separating the A549 and CCL-185-IG groups, while the latter happens as expected for rank 15. This suggests that 15 is a reasonable number of latent factors to include in the model. Importantly, the tested model selection approaches were not informed by cell type labels, which were used only to evaluate the methods’ performance. This shows that unsupervised approaches to model selection are able to estimate a number of factors sufficient to extract meaningful biological signal from real data.
@@
Using 15 factors, as suggested by the model selection criteria, aSGD was compared with NBWaVE, Fisher, and AvaGrad, three of the most popular methods in single-cell analysis, as well as with COAP, one of the fastest and most reliable method in the simulations in addition to aSGD (Supplementary Fig. 3). We tested aSGD with both a Negative Binomial and a Poisson likelihood, and the results were virtually identical (Supplementary Fig. 3). All methods achieve similar results, as shown by both visual inspection of the tSNE plots and the mean cell-line neighborhood purities. However, aSGD is orders of magnitude faster, and therefore allows for model selection obtaining optimal matrix rank, which was shown to be important for clustering the data. Further, aSGD has a better out-of-sample deviance. In terms of memory usage, aSGD has a similar peak RAM memory usage as COAP, outperforming the three other methods.